# Active learning emulators for nuclear two-body scattering in momentum space

### v1

@@ Abstract
We extend the active learning emulators for two-body scattering in coordinate space with error estimation, recently developed by Maldonado et al. in Phys. Rev. C 112, 024002, to coupled-channel scattering in momentum space. Our full-order model (FOM) solver is based on the Lippmann-Schwinger integral equation for the scattering $t$-matrix as opposed to the radial Schrödinger equation. We use (Petrov-)Galerkin projections and high-fidelity calculations at a few snapshots across the parameter space of the interaction to construct efficient reduced-order models (ROMs), trained by a greedy algorithm for locally optimal snapshot selection. Both the FOM solver and the corresponding ROMs are implemented efficiently in Python using Google’s JAX library. We present results for emulating scattering phase shifts in coupled and uncoupled channels and cross sections, and assess the accuracy of the developed ROMs and their computational speed-up factors. We also develop emulator error estimation for both the $t$-matrix and the total cross section. The software framework for reproducing and extending our results will be made publicly available. Together with our recent advances in developing active-learning emulators for three-body scattering, these emulator frameworks set the stage for full Bayesian calibrations of chiral nuclear interactions and optical models against scattering data with quantified emulator errors.
@@ I Introduction
A central challenge in modern nuclear theory is the rigorous quantification of uncertainty, typically achieved through Bayesian statistical methods. These methods require repeated evaluations of high-fidelity models, which can be computationally expensive or even prohibitively slow depending on the application. To overcome this challenge, fast and accurate emulators have recently gained significant popularity in the nuclear physics community. These surrogate models can provide approximate yet highly accurate reduced-order representations of the corresponding high-fidelity (i.e., full-order) models at significantly reduced computational cost. Comprehensive review articles on emulators and their applications to nuclear structure and scattering calculations can be found in Refs. [Melendez:2022kid, Drischler:2022ipa, Duguet:2023wuh].
Among these emulator techniques, the reduced basis method (RBM) has emerged as a powerful and flexible approach. The RBM constructs low-dimensional approximations of the FOM’s solution space using a set of high-fidelity calculations, also called snapshots [hesthaven2015certified, Quarteroni:218966, Benner20201]. These snapshot calculations can be placed naïvely in the parameter space, e.g., using Latin hypercube sampling (LHS) combined with orthonormalization and compression facilitated by the proper orthogonal decomposition (POD), as illustrated in Figure 1 of Ref. [Maldonado:2025ftg]. Alternatively, they can be selected using active learning, applying a so-called greedy algorithm [Sarkar:2021fpz, Bonilla:2022rph, Maldonado:2025ftg] that iteratively places snapshot calculations where the (estimated) emulator error is the largest. It thereby iteratively minimizes the error until a desired accuracy is achieved, preferably with fewer snapshot calculations.
In Ref. [Maldonado:2025ftg], the latter approach was developed and tested for nucleon-nucleon (NN) scattering with chiral interactions in coordinate space, which has been the prototypical test case for emulators in general [Furnstahl:2020abp, Melendez:2021lyq, Drischler:2021qoy, Odell:2023cun, Garcia:2023slj, Catacora-Rios:2025pau]. While fast and accurate emulators for NN scattering are appealing in practice, such emulators are crucial for higher-body scattering, such as nucleon-deuteron (Nd) scattering [Kievsky2008, Marcucci2019, Witala:2021xqm, Gnech:2025gsy, Gnech:2025lbg]. Recently, several of us demonstrated the efficacy of this active learning approach for training proton-deuteron scattering emulators below the deuteron breakup threshold energy [Gnech:2025gsy, Gnech:2025lbg], a crucial step toward full Bayesian calibration of chiral three-nucleon (3N) interactions.
Here, we build on the active-learning emulators for NN scattering in uncoupled partial-wave channels with error estimation, initially developed in Ref. [Maldonado:2025ftg]. We extend them in the following essential ways:
@@
  • high-performance computing (HPC) implementation of these emulators and assessment of their speed-ups.
Together with the 3N emulators developed in Ref. [Gnech:2025gsy, Gnech:2025lbg], this paper provides fast and accurate emulators for calibrating chiral NN and 3N forces efficiently to scattering observables.
To this end, we consider the Lippmann-Schwinger (LS) integral equation in momentum space for the scattering $t$-matrix as our full-order model (FOM). The FOM can be efficiently solved by discretizing the LS equation using an accurate quadrature rule for the principal value integration, resulting in a system of coupled linear equations for the half-on-shell and on-shell matrix elements of the $t$-matrix [Landau:1997]. We then use Galerkin and Petrov-Galerkin projections, along with high-fidelity calculations at a few snapshots of the nuclear interaction parameter space obtained by the greedy algorithm, to construct fast and accurate scattering emulators. Google’s library JAX [jax2018github] enables the efficient implementation of both the FOM solver and the corresponding reduced-order model (ROM) in Python. We demonstrate that Bayesian parameter estimation of nuclear potentials with quantified emulator errors is feasible using our emulators, setting the stage for their applications to rigorous calibrations of chiral nuclear forces. We present results for emulating cross sections and scattering phase shifts in coupled and uncoupled channels using chiral NN interactions and assess the accuracy of the developed ROMs and their computational speed-up factors.
The remainder of this paper is organized as follows. In Sec. II, we describe the formalism of our scattering emulators, which consist of a $t$-matrix high-fidelity solver in momentum space and corresponding ROMs obtained via (Petrov-)Galerkin projections. We then discuss the greedy algorithm in momentum space for the Minnesota potential with a two-dimensional parameter space, which serves as a test case. Next, we present in Sec. III our main results for scattering phase shifts and cross-sections based on more realistic chiral potentials. We also discuss the efficient implementation of our FOM and ROM solvers in Python and discuss the ROM’s speed-up factors. Finally, we perform a proof-of-principle Bayesian parameter calibration of a chiral interaction [Gezerlis:2014zia] to total scattering cross-sections, with emulator errors estimated using the developed methods. This article concludes in Sec. IV with an overview and an outlook. Appendices A and B provide additional information on calculating phase shifts, inelasticity parameters, and total cross sections in momentum space.
We use natural units in which $\hbar=c=1$ and denote matrices and vectors with uppercase and lowercase letters, respectively, both typeset in boldface. The software for reproducing and extending our results will be made publicly available on GitHub [BUQEYEsoftware].
@@ II.1 Full-Order Model: High-Fidelity solution of the NN $t$-matrix equation
where $\widehat{V}$ represents the interaction and $\widehat{G}_{0}(z)=(z-\widehat{H}_{0})^{-1}$, with $z=E+i\varepsilon$ and $\varepsilon>0$ defining the propagator with outgoing boundary conditions. The operator $\widehat{H}_{0}$ stands for the free Hamiltonian, which is given by $\widehat{H}_{0}={\widehat{p}}/2\mu$, with $^{\widehat{p}}$ being the momentum operator and $\mu$ the reduced mass of the NN system. The operator structure of any NN potential is such that only the total angular momentum $j$ is conserved, while the orbital angular momentum $\ell$ is not conserved due to tensor forces. Thus, an angular momentum decomposition of Eq. (1) leads to a single integral equation if $\ell=\ell^{\prime}$,
@@
which simplifies to Eq. (2) without channel coupling (i.e., $\ell=\ell^{\prime}$). The solution and numerical implementation of two-body partial-wave $t$-matrices are well-known in the literature, e.g., as given in Ref. [Glockle:1983]. However, we describe our specific implementation in Appendix A for completeness.
For an efficient offline-online decomposition of our emulator [Drischler:2022ipa, Duguet:2023wuh], we require that the dependence on the parameters $\bf\it\theta$} of the NN potential $V_{\ell,\ell^{\prime}}(k,k^{\prime};{\bf\it\theta})$ is affine so that the potential’s parameter dependence can be separated from its momentum dependence. That means that the partial-wave decomposed potential,
@@
can be written as sums of products of the (smooth) parameter-dependent functions $h_{a}^{(\ell,\ell^{\prime})}({\bf\it\theta})$ and the parameter-independent functions $\mathcal{V}_{a}^{(\ell,\ell^{\prime})}(k,k^{\prime})$. In Eq. (4), we introduced the auxiliary component with constant value ${\bf\it\theta}_{0}\equiv 1$ to accommodate parameter-independent terms in the potential (4). In this work, this extra dimension will encode all pion-exchange contributions to $V_{\ell,\ell^{\prime}}(k,k^{\prime})$, while the other components of the parameter vector $\bf\it\theta$} consist of all short-range low-energy constants (LECs). [^1] In the case of chiral interactions, the parameter dependence on the LECs is even linear; i.e., $h_{a}^{(\ell,\ell^{\prime})}({\bf\it\theta})={\bf\it\theta}_{a}$. We refer to the discussion of the empirical interpolation method (EIM) in, e.g., Refs. [Odell:2023cun, Catacora-Rios:2025pau] and references therein, which can be applied to render potentials with non-affine parameter dependencies approximately affine. Hence, we expect our emulators to cover a wide range of momentum-space potentials, including chiral interactions and optical models. In what follows, we focus on chiral interactions in coupled and uncoupled channels, extending the work in Ref. [Maldonado:2025ftg].
We only need the half-shell partial-wave $t$-matrices to construct the ROMs. Therefore, we can replace the momentum $k^{\prime}$ in Eqs. (2) and (3) by the (fixed) on-shell momentum $k_{0}$, and the single-channel half-shell partial-wave $t$-matrix $T^{j}_{\ell}(k,k_{0};E_{k_{0}})$ becomes a vector of length ($N+1$) when discretized on a momentum grid of size $N$. Here, we choose $N=80$ to obtain high-accuracy results. In the coupled-channel case, the half-shell partial wave $t$-matrix for a given total angular momentum $j$, $T^{j}_{\ell\ell^{\prime}}(k,k_{0};E_{k_{0}})$, with $\ell,\ell^{\prime}=j\mp 1$, will be a vector of length $2(N+1)\times 2(N+1)$ in its discretized form on a momentum grid of size $N$. However, due to the structure of the emulator it is advantageous to rewrite the coupled channel half-shell $t$-matrix as a vector of length $4(N+1)$ as $(T^{j}_{mm},T^{j}_{mp},T^{j}_{pm},T^{j}_{pp})$, where the subscripts indicate the different channels as laid out in Appendix A.
@@
corresponds to the dimensionless scattering amplitude. In terms of phase shifts $\delta^{j}_{\ell\ell^{\prime}}(E_{k_{0}})$, $S^{j}_{\ell\ell^{\prime}}(E_{k_{0}})$ reads [Joachain]
@@
where $\delta^{j}_{\ell\ell^{\prime}}(E_{k_{0}})$ is the scattering phase shift and $\eta^{j}_{\ell\ell^{\prime}}(E_{k_{0}})$ the “inelasticity” or “absorption” factor that has to be one for elastic scattering with real-valued potentials, which is the case considered here. That is, the condition $\eta^{j}_{\ell\ell^{\prime}}(E_{k_{0}})=1$ represents the unitarity of the $S$-matrix. We will use this condition to test if the solutions constructed by the emulator fulfill unitarity. For the explicit calculation of phase shifts and inelasticity parameters from the single- and coupled-channel $S$-matrix elements, we refer to Appendix B, where we follow the Stapp convention [Stapp:1956mz] to obtain the coupled-channel eigen-phase shifts.
@@
Here, we explicitly emphasize the dependence of the matrix elements on the parameters $\bf\it\theta$}. The vector $\mathbf{t}$ contains the half-shell $t$-matrix elements evaluated on the quadrature grid and the on-shell $t$-matrix element. Once the linear system (8) is solved for $\mathbf{t}$ using standard linear algebra methods, we obtain scattering phase shifts at center-of-mass energy $E_{k_{0}}$ in a given partial-wave channel using Eqs. (5)–(7). Furthermore, to compute high-fidelity scattering cross sections, we use the optical theorem, which results in [Silbar:1979wp]
@@ II.2 Reduced-Order Models: (Petrov-)Galerkin projections
Following Ref. [Maldonado:2025ftg], we briefly summarize in this section how ROMs are built using Galerkin projections (G-ROM) and Least-Squares Petrov-Galerkin projections (LSPG-ROM). For a given $\bf\it\theta$}, the FOM of Eq. (8) can be approximated by
@@
In the LSPG-ROM, the coefficient vector $\mathbf{c}({\bf\it\theta})$ is still calculated using Eq. (10). However, instead of forcing the $\mathbf{r}({\bf\it\theta})$ in Eq. (11) to be orthogonal to the reduced space, the coefficient vector $\mathbf{r}({\bf\it\theta})$ is chosen such that the norm of the residual is minimized [Quarteroni:218966], leading to solving the least-squares problem:
@@
Next, $\mathbf{Y}$ is orthonormalized and compressed using a truncated singular value decomposition (SVD), which provides the projection basis. This basis has dimensions $N\times n_{\mathbf{Y}}$, with $n_{\mathbf{Y}}\leqslant(n_{\theta}+1)(n_{b}+1)$. Note that the matrix $\mathbf{Y}\mathbf{Y}^{\dagger}$ is constructed to be an orthogonal projector onto ROM’s approximate subspace of residuals. Finally, this procedure recasts Eq. (13) to obtain the following emulator equation for the LSPG-ROM:
@@
For more details, we refer to Section III.B in Ref. [Maldonado:2025ftg].

### v2

@@ Abstract
We extend the active learning emulators for two-body scattering in coordinate space with error estimation, recently developed by Maldonado et al. [Phys. Rev. C 112, 024002], to coupled-channel scattering in momentum space. Our full-order model (FOM) solver is based on the Lippmann-Schwinger integral equation for the scattering $t$-matrix as opposed to the radial Schrödinger equation. We use (Petrov-)Galerkin projections and high-fidelity calculations at a few snapshots across the parameter space of the interaction to construct efficient reduced-order models (ROMs), trained by a greedy algorithm for locally optimal snapshot selection. Both the FOM solver and the corresponding ROMs are implemented efficiently in Python using Google’s JAX library. We present results for emulating scattering phase shifts in coupled and uncoupled channels and cross sections, and assess the accuracy of the developed ROMs and their computational speedup factors. We also develop emulator error estimation for both the $t$-matrix and the total cross section. The software framework for reproducing and extending our results is publicly available. Together with our recent advances in developing active learning emulators for three-body scattering, these emulator frameworks set the stage for full Bayesian calibrations of chiral nuclear interactions and optical models against scattering data with quantified emulator errors.
@@ I Introduction
A central challenge in modern nuclear theory is the rigorous quantification of uncertainty, typically achieved through Bayesian statistical methods. These methods require repeated evaluations of high-fidelity models, which can be computationally expensive or even prohibitively slow depending on the application. To overcome this challenge, fast and accurate emulators have recently gained significant popularity in the nuclear physics community. These surrogate models can provide approximate yet highly accurate reduced-order representations of the corresponding high-fidelity (i.e., full-order) models at significantly reduced computational cost. Comprehensive review articles on emulators and their applications to nuclear structure and scattering calculations can be found in Refs. [30, 7, 9]. [^1] For original work on projection-based emulators for bound systems, we refer the reader to, e.g., Refs. [14, 10, 26, 6]. A promising alternative approach to accelerating scattering calculations was developed in Ref. [35] using the wave-packet continuum discretization method and GPUs.
Among these emulator techniques, the reduced basis method (RBM) has emerged as a powerful and flexible approach. The RBM constructs low-dimensional approximations of the FOM’s solution space using a set of high-fidelity calculations, also called snapshots [22, 39, 1]. These snapshot calculations can be placed naïvely in the parameter space, e.g., using Latin hypercube sampling (LHS) combined with orthonormalization and compression facilitated by the proper orthogonal decomposition (POD), as illustrated by Maldonado et al. in Figure 1 of Ref. [28]. Alternatively, they can be selected using active learning, applying a so-called greedy algorithm [41, 2, 28] that iteratively places snapshot calculations where the (estimated) emulator error is the largest. It thereby iteratively minimizes the error until a desired accuracy is achieved, preferably with fewer snapshot calculations.
In Ref. [28], the latter approach was developed and tested for nucleon-nucleon (NN) scattering with chiral interactions in coordinate space, which has been the prototypical test case for emulators in general [15, 31, 8, 37, 17, 5]. While fast and accurate emulators for NN scattering are appealing in practice, such emulators are crucial for higher-body scattering, such as nucleon-deuteron (Nd) scattering [25, 29, 50, 20, 21]. Recently, Refs. [20, 21] demonstrated the efficacy of this active learning approach for training proton-deuteron scattering emulators below the deuteron breakup threshold energy, a crucial step toward full Bayesian calibration of chiral three-nucleon (3N) interactions.
Here, we build on the active-learning emulators for NN scattering in uncoupled partial-wave channels with error estimation, initially developed in Ref. [28]. We extend them in the following essential ways:
@@
  • high-performance computing (HPC) implementation of these emulators and assessment of their speedups.
Together with the 3N emulators developed in Ref. [20, 21], this paper provides fast and accurate emulators for calibrating chiral NN and 3N forces efficiently to scattering observables.
To this end, we consider the Lippmann-Schwinger (LS) integral equation in momentum space for the scattering $t$-matrix as our FOM. The FOM can be efficiently solved by discretizing the LS equation using an accurate quadrature rule for the principal value integration, resulting in a system of coupled linear equations for the half-on-shell and on-shell matrix elements of the $t$-matrix [27]. We then use Galerkin and Petrov-Galerkin projections, along with high-fidelity calculations at a few snapshots of the nuclear interaction parameter space obtained by the greedy algorithm, to construct fast and accurate scattering emulators. Google’s library JAX [3] enables the efficient implementation of both the FOM solver and the corresponding reduced-order model (ROM) in Python. We demonstrate that Bayesian parameter estimation of nuclear potentials with quantified emulator errors is feasible using our emulators, setting the stage for their applications to rigorous calibrations of chiral nuclear forces. We present results for emulating cross sections and scattering phase shifts in coupled and uncoupled channels using chiral NN interactions and assess the accuracy of the developed ROMs and their computational speedup factors.
The remainder of this paper is organized as follows. In Sec. II, we describe the formalism of our scattering emulators, which consist of a $t$-matrix high-fidelity solver in momentum space and corresponding ROMs obtained via (Petrov-)Galerkin projections. We then discuss the greedy algorithm in momentum space for the Minnesota potential with a two-dimensional parameter space, which serves as a test case. Next, we present in Sec. III our main results for scattering phase shifts and cross-sections based on more realistic chiral potentials. We also discuss the efficient implementation of our FOM and ROM solvers in Python and discuss the ROM’s speedup factors. Finally, we perform a proof-of-principle Bayesian parameter calibration of a chiral interaction [18] to total scattering cross-sections, with emulator errors estimated using the developed methods. This article concludes in Sec. IV with an overview and an outlook. Appendices A and B provide additional information on calculating phase shifts, inelasticity parameters, and total cross sections in momentum space.
We use natural units in which $\hbar=c=1$ and denote matrices and vectors with uppercase and lowercase letters, respectively, both typeset in boldface. The software for reproducing and extending our results is publicly available on GitHub [4].
@@ II.1 Full-Order Model: High-Fidelity solution of the NN $t$-matrix equation
where $\widehat{V}$ represents the interaction and $\widehat{G}_{0}(z)=(z-\widehat{H}_{0})^{-1}$, with $z=E+i\varepsilon$ and $\varepsilon>0$ defining the propagator with outgoing boundary conditions. The operator $\widehat{H}_{0}$ stands for the free Hamiltonian, which is given by $\widehat{H}_{0}={\widehat{p}}/2\mu$, with $^{\widehat{p}}$ being the momentum operator and $\mu$ the reduced mass of the NN system. The operator structure of any NN potential is such that while the total angular momentum $j$ is conserved, the orbital angular momentum $\ell$ is not conserved due to tensor forces. Thus, an angular momentum decomposition of Eq. (1) leads to a single integral equation if $\ell=\ell^{\prime}$,
@@
which simplifies to Eq. (2) without channel coupling (i.e., $\ell=\ell^{\prime}$). The solution and numerical implementation of two-body partial-wave $t$-matrices are well-known in the literature, e.g., as given in Ref. [19]. However, we describe our specific implementation in Appendix A for completeness.
For an efficient offline-online decomposition of our emulator [7, 9], we require that the dependence on the parameters $\bf\it\theta$} of the NN potential $V_{\ell,\ell^{\prime}}(k,k^{\prime};{\bf\it\theta})$ is affine so that the potential’s parameter dependence can be separated from its momentum dependence. That means that the partial-wave decomposed potential,
@@
can be written as sums of products of the (smooth) parameter-dependent functions $h_{a}^{(\ell,\ell^{\prime})}({\bf\it\theta})$ and the parameter-independent functions $\mathcal{V}_{a}^{(\ell,\ell^{\prime})}(k,k^{\prime})$. In Eq. (4), we introduced the auxiliary component with constant value ${\bf\it\theta}_{0}\equiv 1$ to accommodate parameter-independent terms in the potential (4). In this work, this extra dimension will encode all pion-exchange contributions to $V_{\ell,\ell^{\prime}}(k,k^{\prime})$, while the other components of the parameter vector $\bf\it\theta$} consist of all short-range low-energy constants (LECs). [^2] In the case of chiral interactions, the parameter dependence on the LECs is even linear; i.e., $h_{a}^{(\ell,\ell^{\prime})}({\bf\it\theta})={\bf\it\theta}_{a}$. We refer to the discussion of the empirical interpolation method (EIM) in, e.g., Refs. [37, 5] and references therein, which can be applied to render potentials with non-affine parameter dependencies approximately affine. Hence, we expect our emulators to cover a wide range of momentum-space potentials, including chiral interactions and optical models. In what follows, we focus on chiral interactions in coupled and uncoupled channels.
We only need the half-shell partial-wave $t$-matrices to construct the ROMs. Therefore, we can replace the momentum $k^{\prime}$ in Eqs. (2) and (3) by the (fixed) on-shell momentum $k_{0}$, and the single-channel half-shell partial-wave $t$-matrix $T^{j}_{\ell}(k,k_{0};E_{k_{0}})$ becomes a vector of length ($N+1$) when discretized on a momentum grid of size $N$. Here, we choose $N=80$ to obtain high-accuracy results and truncate the numerical integration at a sufficiently high-enough maximum momentum, $k_{\mathrm{max}}=20\,\text{fm}^{-1}$. In the coupled-channel case, the half-shell partial wave $t$-matrix for a given total angular momentum $j$, $T^{j}_{\ell\ell^{\prime}}(k,k_{0};E_{k_{0}})$, with $\ell,\ell^{\prime}=j\mp 1$, will be a vector of length $2(N+1)\times 2(N+1)$ in its discretized form on a momentum grid of size $N$. However, due to the structure of the emulator it is advantageous to rewrite the coupled channel half-shell $t$-matrix as a vector of length $4(N+1)$ as $(T^{j}_{mm},T^{j}_{mp},T^{j}_{pm},T^{j}_{pp})$, where the subscripts indicate the different channels as laid out in Appendix A.
@@
corresponds to the dimensionless scattering amplitude. In terms of phase shifts $\delta^{j}_{\ell\ell^{\prime}}(E_{k_{0}})$, $S^{j}_{\ell\ell^{\prime}}(E_{k_{0}})$ reads [24]
@@
where $\delta^{j}_{\ell\ell^{\prime}}(E_{k_{0}})$ is the scattering phase shift and $\eta^{j}_{\ell\ell^{\prime}}(E_{k_{0}})$ the “inelasticity” or “absorption” factor that has to be one for elastic scattering with real-valued potentials, which is the case considered here. That is, the condition $\eta^{j}_{\ell\ell^{\prime}}(E_{k_{0}})=1$ represents the unitarity of the $S$-matrix. We will use this condition to test if the solutions constructed by the emulator fulfill unitarity. For the explicit calculation of phase shifts and inelasticity parameters from the single- and coupled-channel $S$-matrix elements, we refer to Appendix B, where we follow the Stapp convention [43] to obtain the coupled-channel eigen-phase shifts.
@@
Here, we explicitly emphasize the dependence of the matrix elements on the parameters $\bf\it\theta$}. The vector $\mathbf{t}$ contains the half-shell $t$-matrix elements evaluated on the quadrature grid and the on-shell $t$-matrix element. Once the linear system (8) is solved for $\mathbf{t}$ using standard linear algebra methods, we obtain scattering phase shifts at center-of-mass energy $E_{k_{0}}$ in a given partial-wave channel using Eqs. (5)–(7). Furthermore, to compute high-fidelity scattering cross sections, we use the optical theorem, which results in [42]
@@ II.2 Reduced-Order Models: (Petrov-)Galerkin projections
We briefly summarize in this section how ROMs are built using Galerkin projections (G-ROM) and Least-Squares Petrov-Galerkin projections (LSPG-ROM). For a given $\bf\it\theta$}, the FOM of Eq. (8) can be approximated by
@@
In the LSPG-ROM, the coefficient vector $\mathbf{c}({\bf\it\theta})$ is still calculated using Eq. (10). However, instead of forcing the $\mathbf{r}({\bf\it\theta})$ in Eq. (11) to be orthogonal to the reduced space, the coefficient vector $\mathbf{r}({\bf\it\theta})$ is chosen such that the norm of the residual is minimized [39], leading to solving the least-squares problem:
@@
Next, $\mathbf{Y}$ is orthonormalized and compressed using a truncated singular value decomposition (SVD), which provides the projection basis. This basis has dimensions $N\times n_{\mathbf{Y}}$, with $n_{\mathbf{Y}}\leqslant(n_{\theta}+1)(n_{b}+1)$. Note that the matrix $\mathbf{Y}\mathbf{Y}^{\dagger}$ is constructed to be an orthogonal projector onto the ROM’s approximate subspace of residuals. Finally, this procedure recasts Eq. (13) to obtain the following emulator equation for the LSPG-ROM:
@@
For more details, we refer to Section III.B in Ref. [28].